# Large Language Models with Task-Oriented Experts

### v1

@@ 4.3 Post-recognition LLMs Setup
We implement NeKo using the Transformer architecture (Vaswani et al., 2017) and fine-tune both dense and MoE models for comparison. For dense models, we fine-tune Gemma 2B (Team et al., 2024) and Mistral 7B (Jiang et al., 2024b). For MoE models, we fine-tune Gemma 8x2B [^2] We made an up-cycled (Komatsuzaki et al., 2023) Gemma 8x2B MoE setup extended from single Gemma-2B (Team et al., 2024). and Mixtral 8x7B without applying our task-oriented expert assignment. We explore the Branch-Train-Mix approach (Sukhbaatar et al., 2024), which involves branching from the Mistral 7B model to an 8x7B MoE model as one competing setup. To investigate the scalability of our method, we design NeKo to three different sizes of MoE models: Gemma 8x2B, Mixtral 8x7B, and Mixtral 8x22B. We further compared low-rank adaptation (LoRA(Hu et al., 2021)) with full fine-tuning (FFT) on 8x7B MoE setup.
For MoE models, we use top-k routing as proposed in (Lepikhin et al., 2021) to balance the computational cost and model capacity. We use a global batch size of 2 million tokens and apply sample packing (Raffel et al., 2020) to maximize the GPU utilization. We fine-tune the model for 3 epochs using the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of $41e-$ 4 and a weight decay of $0.01$. We use a cosine learning rate scheduler with a warmup ratio of $0.1$ and a gradient clipping threshold of $1.0$. For the expert-dataset mapping, we randomly assign each dataset to one of the 8 experts in the Mixtral model. This random assignment serves as a strong baseline and allows us to focus on the effectiveness of the task-oriented expert assignment approach. We leave the exploration of more advanced expert assignment strategies for future work. To efficiently train the large-scale model, we leverage DeepSpeed Zero (Rajbhandari et al., 2020) for memory optimization and Hugging Face Transformers (Wolf et al., 2020) for model implementation.
@@
NeKo-FT 26.9 44.2 26.3 27.7 44.4 36.0 34.3 20.7 29.6 43.4 31.2 33.1 28.2 16.8 26.0
NeKo-BTX 27.2 44.5 26.7 28.0 44.7 36.3 34.6 21.0 29.9 43.8 31.6 33.4 28.5 17.1 26.3
NeKo-MoE 28.5 46.2 28.0 30.1 46.3 38.7 36.3 23.4 32.6 46.5 34.2 37.2 32.8 21.5 30.5
4.4 Post-recognition Generative Correction Results
@@
We first evaluate the zero-shot ability of NeKo on unseen domain compared to two general-purpose LLMs, including GPT-3.5 Turbo and Claude-Opus [^3] Claude-Opus is one of the high-performance, general-purpose LLMs evaluated in the LLM evaluation (Zheng et al., 2024) and performs competitively compared to GPT-4-Turbo and Gemini 1.5-Pro, offering unlimited API access.. With a task-specific recognition baseline of Whisper-V2-Large (third column) in Table 1, NeKo-MoE shows the best zero-shot ability with a relative 22.3% average WER reduction. GPT-3.5 Turbo and Claude-Opus have relative 4.3% and 7.3% of zero-shot improvements, where NeKo consistently outperform their 5-shot ASR correction.
@@
Compared to other leading models on the leaderboard, NeKo establishes a new state-of-the-art, outperforming speech-only foundational models like Whisper and Canary and end-to-end ASR-LLM like SALM  [^4] We conduct one competitive end-to-end, multi-modal baseline based on Bestow (Chen et al., 2024c) with joint audio and text injection under cross-attention, which shows improved performance on the Open ASR benchmark. (Chen et al., 2024b) across most datasets. On the AMI dataset, NeKo achieves a WER of 12.58%, significantly lower than Whisper’s 16.82%. On VoxPopuli, NeKo obtains 5.84% WER, a 1.66 point reduction from Whisper’s 7.5%. The strong performance of NeKo demonstrates the effectiveness of our speech-adapted MoE approach in handling diverse speech datasets and learning robust representations.
@@
Table 3 presents the speech translation results on the FLEURS, CoVoST-2, and MuST-C datasets. For these experiments, we use SeamlessM4T-Large as the task-specific model to generate the initial speech translation hypotheses. NeKo is then applied to correct the outputs from SeamlessM4T-Large. Compared to the task-specific SeamlessM4T-Large model, NeKo achieves significant improvements, with an average BLEU score increase of 5.4 points on the FLEURS dataset, 9.2 points on the CoVoST-2 dataset, and 5.4 points on the MuST-C dataset. These results demonstrate the effectiveness of NeKo in correcting errors made by the first-pass speech translation model. Moreover, NeKo outperforms other correction baselines, including the state-of-the-art GenTranslate model. On the FLEURS dataset, NeKo obtains an average BLEU score of 36.3, surpassing GenTranslate by 2.3 points. For CoVoST-2, NeKo achieves an average BLEU score of 34.2, outperforming GenTranslate by 3.2 points. On MuST-C, NeKo improves over GenTranslate by 3.2 BLEU points, reaching an average score of 30.5.
Table 4: Machine translation BLEU scores on the WMT’20 Japanese (Ja) and Chinese (Zh) test sets (Barrault et al., 2020b). NeKo is evaluated in a zero-shot setting, while other models are fine-tuned on the respective language pairs. Higher BLEU scores indicate better translation quality.
@@
NeKo-MoE (zero-shot) 18.1 27.6 22.9
To further assess the generalization ability of NeKo , we evaluate it on the WMT’20 machine translation benchmark for Japanese and Chinese in a zero-shot setting. As shown in Table 4, NeKo achieves competitive performance compared to fine-tuned MT models, obtaining an average BLEU score of 22.9. This result highlights the potential of NeKo to handle unseen translation tasks by leveraging the knowledge learned from speech translation datasets.
@@
For the OCR task, NeKo achieves a substantial error reduction, lowering the WER from 71.03% to 14.43%. This represents a significant improvement over the baseline and demonstrates the model’s ability to correct OCR errors effectively. Compared to the Mixtral-MoE model fine-tuned directly on the OCR dataset, NeKo obtains a 1.02% lower WER, highlighting the benefit of the task-oriented expert assignment approach. In the TEC task, NeKo showcases its versatility by improving the performance on both grammar correction and coherence improvement subtasks. For grammar correction, NeKo reduces the WER from 31.41% to 9.42%, outperforming the directly fine-tuned Mixtral-MoE model by 1.31%. On the coherence subtask, NeKo achieves a WER of 9.71%, which is 0.46% higher than the directly fine-tuned model but still a significant improvement over the baseline.
Table 5: WER comparison of NeKo against the baseline and a directly fine-tuned Mixtral-MoE model (8x7B) on grammar correction and coherence improvement tasks from the CoEdIT dataset (Raheja et al., 2023), and the OCR task using the PleIAs/Post-OCR-Correction dataset (PleIAs, 2023). Lower WER indicates better performance.
Task Grammar Correction Coherence Improv. OCR
Baseline 31.41 13.48 71.03
Mixtral-MoE-FFT 10.73   9.25 15.45
NeKo-MoE   9.42   9.71 14.43
Model WER
Baseline (Corrupt vs Ground Truth) 18.35%
GPT-3.5-Turbo (5-shots) 12.72%
Claude-Opus (5-shots) 8.18%
NeKo-MoE (5-shots) 11.62%
Emergent Unseen Task Zero-Shot Performance
@@ 5 Conclusion
In this work, we proposed NeKo, a multi-task GER approach that leverages task-oriented MoEs to handle diverse tasks. NeKo assigns each expert to a specific dataset during training, enabling the experts to capture task-specific features while allowing knowledge sharing through the gating network. Our results show that task-oriented expert assignment is a promising approach for multi-task learning in error correction and other natural language processing tasks. By aligning experts with datasets, NeKo can effectively capture the nuances and specificities of each task while benefiting from the shared knowledge learned by the gating network and other model components. Future work includes exploring more advanced expert assignment strategies, such as dynamically assigning experts based on the input characteristics. Investigating the interpretability of the learned expert representations and routing decisions would be important next step upon the open NeKo models.
Limitation
@@
Task-Specific Fine-Tuning
The NeKo model employs task-oriented MoE fine-tuning, where each expert is assigned to a specific dataset. This approach may lead to overfitting to the specific characteristics of the training datasets even though knowledge could be shared. As a result, the model’s performance might degrade when applied to new tasks or datasets that were not part of the training set, limiting its adaptability. Investigating more dynamic and adaptive fine-tuning strategies that can generalize better across unseen tasks and datasets would be beneficial. Techniques such as meta-learning or continual learning could be explored to enhance the model’s adaptability and robustness.
Future Connections to In-Context and Auto-Agent Learning with NeKo
@@ Appendix A Appendix
Figure 6: Examples of NeKo outputs for machine translation correction task in WMT20 Barrault et al. (2020a).
@@
Task-Oriented Inference for Mixture of Expert Models
@@
Using more than two experts (e.g., top-3 or higher) diverged from the training setup and increased inference costs (ranging from $23.5$% to $75.5$%) without significant gain (i.e., a relative difference of less than $0.06$%).
@@ Acknowledgment
The authors would like to acknowledge the initial feedback from Taejin Park, Boyi Li, and Yejin Choi at NVIDIA and the discussion on medical speech recognition evaluation with Zhen Wan and Chenhui Chu from Kyoto University.

### v2

@@ 3.3 Post-recognition LLMs Setup
We implement NeKo using the Transformer architecture (Vaswani et al., 2017) and fine-tune both dense and MoE models for comparison. For dense models, we fine-tune Gemma 2B (Team et al., 2024) and Mistral 7B (Jiang et al., 2024b). For MoE models, we fine-tune Gemma 8x2B [^1] We made an up-cycled (Komatsuzaki et al., 2023) Gemma 8x2B MoE setup extended from single Gemma-2B (Team et al., 2024). and Mixtral 8x7B without applying our task-guided expert assignment. We explore the Branch-Train-Mix approach (Sukhbaatar et al., 2024), which involves branching from the Mistral 7B model to an 8x7B MoE model as one competing setup. To investigate the scalability of our method, we design NeKo to three different sizes of MoE models: Gemma 8x2B, Mixtral 8x7B, and Mixtral 8x22B. We further compared low-rank adaptation (LoRA(Hu et al., 2021)) with full fine-tuning (FFT) on 8x7B MoE setup.
For MoE models, we use top-k routing as proposed in (Lepikhin et al., 2021) to balance the computational cost and model capacity. We use a global batch size of 2 million tokens and apply sample packing (Raffel et al., 2020) to maximize the GPU utilization.
@@
NeKo-Gemma-2B-FT 26.9 44.2 26.3 27.7 44.4 36.0 34.3 20.7 29.6 43.4 31.2 33.1 28.2 16.8 26.0
NeKo-Gemma-8x2B–BTX 27.2 44.5 26.7 28.0 44.7 36.3 34.6 21.0 29.9 43.8 31.6 33.4 28.5 17.1 26.3
NeKo-Gemma-8x2B-MoE 28.5 46.2 28.0 30.1 46.3 38.7 36.3 23.4 32.6 46.5 34.2 37.2 32.8 21.5 30.5
3.4 Post-recognition Correction Results
@@
We first evaluate the zero-shot ability of NeKo on unseen domain compared to two general-purpose LLMs, including GPT-3.5 Turbo and Claude-3.5 Sonnet. With a task-specific recognition baseline of Whisper-V2-Large (third column) in Table 1, NeKo-MoE (i.e., Qwen1.5-MoE or Mixtral) shows the best zero-shot ability with a relative 22.3% average WER reduction. GPT-3.5 Turbo and Claude-3.5 Sonnet have relative 4.3% and 7.3% of zero-shot improvements, where NeKo consistently outperform their 5-shot ASR correction.
@@
Compared to other leading models on the leaderboard, NeKo establishes a new state-of-the-art, outperforming speech-only foundational models like Whisper and Canary and end-to-end ASR-LLM like SALM (Chen et al., 2024b) across most datasets. On the AMI dataset, NeKo achieves a WER of 12.58%, significantly lower than Whisper’s 16.82%. On VoxPopuli, NeKo obtains 5.84% WER, a 1.66 point reduction from Whisper’s 7.5%. The strong performance of NeKo demonstrates the effectiveness of our speech-adapted MoE approach in handling diverse speech datasets and learning robust representations.
@@
Table 3 presents the speech translation results on the FLEURS, CoVoST-2, and MuST-C datasets. For these experiments, we use SeamlessM4T-Large as the task-specific model to generate the initial speech translation hypotheses. NeKo is then applied to correct the outputs from SeamlessM4T-Large. Compared to the task-specific SeamlessM4T-Large model, NeKo achieves significant improvements, with an average BLEU score increase of 5.4 points on the FLEURS dataset, 9.2 points on the CoVoST-2 dataset, and 5.4 points on the MuST-C dataset. These results demonstrate the effectiveness of NeKo in correcting errors made by the first-pass speech translation model. Moreover, NeKo outperforms other correction baselines, including the state-of-the-art GenTranslate model.
Table 4: Machine translation BLEU scores on the WMT’20 Japanese (Ja) and Chinese (Zh) test sets (Barrault et al., 2020a). NeKo is evaluated in a zero-shot setting, while other models are fine-tuned on the respective language pairs. Higher BLEU scores indicate better translation quality.
@@
NeKo-Gemma-MoE (zero-shot) 18.1 27.6 22.9
To further assess the generalization ability of NeKo , we evaluate it on the WMT’20 machine translation benchmark for Japanese and Chinese in a zero-shot setting. As shown in Table 4, NeKo achieves competitive performance compared to fine-tuned MT models, obtaining an average BLEU score of 22.9. This result highlights the potential of NeKo to handle unseen translation tasks by leveraging the knowledge learned from pre-training.
@@
For the OCR task, NeKo achieves a substantial error reduction, lowering the WER from 71.03% to 14.43%. This represents a significant improvement over the baseline and demonstrates the model’s ability to correct OCR errors effectively. Compared to the Mixtral-MoE model fine-tuned directly on the OCR dataset, NeKo obtains a 1.02% lower WER, highlighting the benefit of the task-guided expert assignment approach. In the TEC task, NeKo showcases its versatility by improving the performance on both grammar correction and coherence improvement subtasks. For grammar correction, NeKo reduces the WER from 31.41% to 9.42%, outperforming the directly fine-tuned Mixtral-MoE model by 1.31%. On the coherence subtask, NeKo achieves a WER of 9.71%, which is 0.46% higher than the directly fine-tuned model but still a significant improvement over the baseline.
Table 5: WER comparison of NeKo against the baseline and a directly fine-tuned Mixtral-MoE model (8x7B) on grammar correction and coherence improvement tasks from the CoEdIT dataset (Raheja et al., 2023), and the OCR task using the PleIAs/Post-OCR-Correction dataset (PleIAs, 2023).
Task / WER $\downarrow$ Grammar Correction Coherence Improv. OCR
Mixtral-MoE (frozen) 31.41 13.48 71.03
GPT-3.5-turbo 17.43   12.25 39.45
Mixtral-MoE-FFT 10.73   12.05 45.32
NeKo-Mixtral-MoE   9.42   9.71 14.43
@@ 4 Conclusion
In this work, we proposed NeKo, a multi-task GER approach that leverages task-guided MoEs to handle diverse tasks. NeKo assigns each expert to a specific dataset during training, enabling the experts to capture task-specific features while allowing knowledge sharing through the gating network. Our results show that task-guided expert assignment is a promising approach for multi-task learning in error correction and other natural language processing tasks. By aligning experts with datasets, NeKo can effectively capture the nuances and specificities of each task while benefiting from the shared knowledge learned by the gating network and other model components. Future work includes exploring more advanced expert assignment strategies, such as dynamically assigning experts based on the input characteristics.
Ethical Considerations
@@ Appendix A Appendix
Training Details
We fine-tune the model for 3 epochs using the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of 1e41e-4 and a weight decay of $0.01$. We use a cosine learning rate scheduler with a warmup ratio of $0.1$ and a gradient clipping threshold of $1.0$. For the expert-dataset mapping, we randomly assign each dataset to one of the 8 experts in the Mixtral model. This random assignment serves as a strong baseline and allows us to focus on the effectiveness of the task-guided expert assignment approach. We leave the exploration of more advanced expert assignment strategies for future work. To efficiently train the large-scale model, we leverage DeepSpeed Zero (Rajbhandari et al., 2020) for memory optimization and Hugging Face Transformers (Wolf et al., 2020) for model implementation.
Translation Tasks. As an extra zero-shot textual correction setup, we evaluate NeKo on machine translation (MT) of WMT’20 for Japanese and Chinese (Barrault et al., 2020b). We use the BLEU score (Papineni et al., 2002) as the evaluation metric for ST (with training and test) and MT (zero-shot).
Grammar Correction Tasks. These text error correction (TEC) tasks focus on correcting grammatical errors and improving the overall coherence of the text, making them suitable for evaluating the effectiveness of our model in handling TEC-related editing instructions. We use the word error rate as the evaluation metric.
OCR Tasks. The dataset includes original texts with varying numbers of OCR mistakes and their corresponding corrected versions. To evaluate our model, we take the first 1,000 characters of both the input text with OCR errors and the ground-truth corrected text. We use the WER as the evaluation metric.
Mixture of Experts Background
Mixture-of-experts (MoE) (Shazeer et al., 2017) is a machine learning concept that employs multiple expert layers, each of which specializes in solving a specific subtask. The experts then work together to solve the entire task at hand. Recently, MoE has been widely applied to large-scale distributed Deep Learning models by using a cross-GPU layer that exchanges hidden features from different GPUs (Lepikhin et al., 2021; Fedus et al., 2022). The MoE approach is differentiated from existing scale-up approaches for DNNs, such as increasing the depth or width of DNNs, in terms of its high cost-efficiency. Specifically, adding more model parameters (experts) in MoE layers does not increase the computational cost per token at inference time. Thus, MoE has been studied for scaling the models to trillion-size parameters in NLP (Fedus et al., 2022).
@@
Figure 6: Examples of NeKo outputs for machine translation correction task in WMT20 Barrault et al. (2020b).
@@
task-guided Inference for Mixture of Expert Models
@@
Using more than two experts (e.g., top-3 or higher) diverged from the training setup and increased inference costs (ranging from 23.523.5% to 75.575.5%) without significant gain (i.e., a relative difference of less than 0.060.06%).
@@
Emergent Unseen Task Zero-Shot Performance
Task-Specific Fine-Tuning
The NeKo model employs task-guided MoE fine-tuning, where each expert is assigned to a specific dataset. This approach may lead to overfitting to the specific characteristics of the training datasets even though knowledge could be shared. As a result, the model’s performance might degrade when applied to new tasks or datasets that were not part of the training set, limiting its adaptability. Investigating more dynamic and adaptive fine-tuning strategies that can generalize better across unseen tasks and datasets would be beneficial. Techniques such as meta-learning or continual learning could be explored to enhance the model’s adaptability and robustness.
Future Connections to In-Context and Auto-Agent Learning with NeKo
Model WER
Baseline (Corrupt vs Ground Truth) 18.35%
GPT-3.5-Turbo (5-shots) 12.72%
Claude-3-Sonnet (5-shots) 12.18%
Claude-3.5 Sonnet (5-shots) 8.18%
NeKo-MoE (5-shots) 11.62%
@@ Acknowledgment
The authors would like to acknowledge the Taipei-1 high-performance computing support provided by the NVIDIA Taiwan Research and Development Center (TRDC) project, and the initial feedback from NVIDIA’s Eric Kang, Taejin Park, Boyi Li, and Yejin Choi.